# The Feldman-Hájek Dichotomy for Countable Gaussian Mixtures and their Asymptotic Separability in High Dimensions

### v1

@@
Abstract
This paper establishes the theoretical foundations for the asymptotic separability of Gaussian Mixture Models (GMMs) in high dimensions by extending the classical Feldman-Hájek theorem. We first prove that a countable mixture of Gaussian measures is a well-defined probability measure. Our primary result, the Gaussian Mixture Dichotomy Theorem, demonstrates that the mutual singularity of individual Gaussian components is a sufficient condition for the mutual singularity of the resulting mixtures. We provide a rigorous proof and further discuss the “Mixed Case,” where the presence of even a single equivalent pair of components leads to partial absolute continuity via the Lebesgue decomposition, thereby defining the theoretical limits of perfect classification in infinite-dimensional spaces.
Infinite-Dimensional Measures,
keywords:
[class=MSC]
keywords:
@@ 1 Introduction
The study of probability measures in high-dimensional spaces is often dominated by the phenomenon of measure concentration [2, 3, 4, 5]. As $d\to\infty$ ($d$ being the dimension), the geometry of these measures changes radically, leading to the behavior described for Gaussian measures in 1958 by the Feldman-Hájek dichotomy [4, 5]. This result belongs to a broader class of dichotomy theorems for infinite-dimensional measures, the most notable being Kakutani’s theorem on infinite product measures [7]. The Feldman-Hajék theorem states that two Gaussian measures in an infinite-dimensional Hilbert space are either equivalent or mutually singular. While the behavior of individual Gaussian measures is well-documented, modern statistical learning frequently employs Gaussian Mixture Models (GMMs) to represent complex distributions, and thus it is a relevant questions under which conditions two different GMMs can be perfectly separable in infinite dimensions. Some general discussion about sums of random variables and mixtures can be found in the book by Vakhania et al. in [12], but not a specific result for GMMs.
@@
Several authors have studied GMM models in high dimensional settings, but nobody, to the best knowledge of the author, has given conditions for mutual singularity of GMM models in infinite dimensions. Almost all research work found, dealt with the problem of classifying or learning GMM models, and were not interested in finding conditions for mutually singularity.
@@
It is maybe of interest to the reader to highlight some of the works related to learning GMMs. It should be emphasized that these contributions are, in a certain sense, connected to and conceptually aligned with the research presented in this work. In fact, if two GMMs are mutually singular, it also means that they are perfectly separable, or in other words, in the assumption that one finds the right classifier, one can classify them with zero missclassification error. A first interesting work in this direction, is the one by Mignacco et al. [9]. In their work, they studied the role of regularisation in a classification problem of high-dimensional two-mixture of noisy gaussians. Refinetti et al. studied the problem of classifying high-dimensional Gaussian mixtures, with a focus on neural networks [10] and showed the superiority of neural networks over kernel methods. Azizyan et al. in [1] also studied the problem of learning Gaussian mixtures with equal mixing weights in a high-dimensional settings under small-mean separation. Other works goes along the same line, in studying how to learn gaussian mixtures. The objective of this study is not to address that point directly, but rather to investigate how the Feldman–Hajék theorem can be generalized and extended to the setting of Gaussian mixture measures. As already mentioned, this work is important for the problem of learning mixtures, as it deals with the limit where error-free (perfect) classification becomes theoretically possible.
@@
The primary contribution of this work is the extension of the Feldman-Hájek theorem to countable Gaussian Mixture Models. Specifically, we establish the following. (i) We provide a rigorous proof that a countable mixture of Gaussian measures is a well-defined probability measure on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$. (ii) We prove that the mutual singularity of all constituent Gaussian components is a sufficient condition for the mutual singularity of the resulting mixtures. (iii) We discuss shortly the “mixed case,” demonstrating how the existence of equivalent components breaks the dichotomy and leads to partial absolute continuity. By extending these results to GMMs, we provide a theoretical foundation for the study of asymptotic separability in complex, multi-modal distributions.
@@
The rest of this paper is organised as follows. In Section 2, we establish the formal measure-theoretic framework for countable mixtures and prove that they are valid probability measures via Lemmas 2.1 and 2.2. Section 3 presents our main result, the The Gaussian Mixture Dichotomy Theorem (Theorem 3.1), and its proof. Finally, Section 4 discusses the implications of these results.
@@ 2 Formalism and Gaussian Mixtures
Let $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ be a measurable space, where $\mathcal{B}(\mathbb{R}^{d})$ is the Borel $\sigma$-algebra. A countable mixture is a probability measure $\mu$ defined for every $A\subseteq\mathcal{B}(\mathbb{R}^{d})$ as the following.
@@
where $\gamma_{k}$ are probability Gaussian measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ characterised by the mean vector $m_{k}\in\mathbb{R}^{d}$ and by their covariance $\Sigma_{k}$. $\pi_{k}$ are mixing weights such that $\pi_{k}>0$ and $\sum_{k=1}^{\infty}\pi_{k}=1$. In Lemma 2.2 we prove that $\mu$ itself is a probability measure. Lemma 2.1 is a stepping stone to prove lemma 2.2.
@@
Given a Gaussian measure $\gamma_{i}$, for any For any sequence of disjoint sets $\{A_{n}\}_{n=1}^{\infty}$ with $A_{n}\subset\mathcal{B}(\mathbb{R}^{d})$ the sum $\sum_{k=1}^{\infty}\gamma_{i}(A_{k})$ exists and is finite.
Proof.
@@
But also by definition $\cup_{n=1}^{\infty}A_{n}\subseteq\mathbb{R}^{d}$ and therefore
$\gamma_{i}\left(\bigcup_{n=1}^{\infty}A_{n}\right)\leq\gamma_{i}(\mathbb{R}^{d})=1$ (3)
Additionally, since each $\gamma_{i}(A_{n})\geq 0$ then the partial sums $s_{m}=\sum_{n=1}^{m}\gamma_{i}(A_{n})$ are monotonically increasing and bounded above by 1, hence it converges. ∎
Lemma 2.2 ($\mu$ is a measure).
@@
where $\gamma_{k}$ are probability Gaussian measures on $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ characterised by the mean vector $m_{k}\in\mathbb{R}^{d}$ and by their covariance $\Sigma_{k}$. $\pi_{k}$ are mixing weights such that $\pi_{k}>0$ and $\sum_{k=1}^{\infty}\pi_{k}=1$. $\mu$ is a probability measure.
Proof.
@@
Now let us consider two GMMs $\mu$ and $\nu$ defined over two disjoint sets of indices $I_{1}$ and $I_{2}$ similarly to Equation 1. Namely we will define
@@
Now we will prove the following lemma.
Lemma 2.3.
If $\gamma_{i}\perp\gamma_{j}$ (the two measures are mutually singular) $\forall i\in I_{1}$ and $\forall j\in I_{2}$ then $\mu\perp\nu$.
Proof.
To prove this lemma, we need to show that $\exists A$ such that $\mu(A)=1$ and $\nu(A)=0$. Let us construct such a set. Recall that, by assumption, $\gamma_{i}\perp\gamma_{j}$ (the two measures are mutually singular) $\forall i\in I_{1}$ and $\forall j\in I_{2}$, thus there exists a set $A_{ij}$ such that $\gamma_{i}(A_{ij})=1$ and $\gamma_{j}(A_{ij})=0$. Consider the set
@@
$\bigcup_{i\in I_{1}}S_{i}\subseteq\mathbb{R}^{d}\Rightarrow\mu\left(\bigcup_{i\in I_{1}}S_{i}\right)\leq\mu(\mathbb{R}^{d})=1$ (18)
@@
Or in other words we have proven that $1\leq\mu(A)\leq 1$ and therefore $\mu(A)=1$. Thus we have proven that $\mu\perp\nu$. ∎
@@ 3 The Gaussian Mixture Dichotomy Theorem
Let $(\mathbb{R}^{d},\mathcal{B}(\mathbb{R}^{d}))$ be a measurable space. Let $\mu$ and $\nu$ be two Gaussian Mixture Models (GMMs) defined over disjoint index sets $I_{1}$ and $I_{2}$ as
@@
    The mean difference $(m_{i}-m_{j})$ does not belong to the common Cameron-Martin Space $H=\Sigma_{i}^{1/2}(\mathbb{R}^{d})=\Sigma_{j}^{1/2}(\mathbb{R}^{d})$.
Proof.
By Lemma 2.3, if $\gamma_{i}\perp\gamma_{j}$ for all $i\in I_{1}$ and $j\in I_{2}$, then $\mu\perp\nu$. Under the standard Feldman-Hájek theorem, two Gaussian measures are mutually singular if and only if the aforementioned operator or mean conditions are violated. Since these conditions imply $\gamma_{i}\perp\gamma_{j}$, the singularity of the mixtures follows from the existence of the set $A=\bigcup_{i\in I_{1}}(\bigcap_{j\in I_{2}}A_{ij})$ such that $\mu(A)=1$ and $\nu(A)=0$. ∎
@@ 4 Results and Discussion
In finite dimensions ($\mathbb{R}^{d}$), two non-degenerate Gaussian measures are almost always equivalent because their supports coincide with the entire space. However, singularity ($\gamma_{i}\perp\gamma_{j}$) can occur in finite dimensions if the measures are degenerate, for instance, if they are supported on disjoint lower-dimensional affine subspaces.
The The Gaussian Mixture Dichotomy Theorem becomes interesting as $d\to\infty$. In infinite-dimensional spaces, the dichotomy takes hold: two Gaussian measures are either equivalent or mutually singular. This rigorous disjointness of support in infinite dimensions is a characteristic property of Gaussian measures [4] and is in general valid for product-type measures [7] and, as we have demonstrated, scales to Gaussian Mixture Models (GMMs). If the components of the mixtures are singular, a condition almost always satisfied in the infinite-dimensional limit due to shifts in mean or scaling of covariance, the global mixtures $\mu$ and $\nu$ are themselves perfectly separable.
@@ 4.2 The Mixed Case
An important aspect in applying the The Gaussian Mixture Dichotomy Theorem arises when the index sets $I_{1}$ and $I_{2}$ each include Gaussian pairs that are not all of the same type, but instead comprise a combination of pairs that are equivalent and pairs that are mutually singular. While the standard theorem for pure Gaussian measures permits only a binary state (equivalence or singularity), the mixture model $\mu$ and $\nu$ allows for a ”mixed” or intermediate state.
@@
In practical terms, this means that the strictly “all-or-nothing” nature of separability in GMMs fundamentally depends on the pair-wise singularity condition being satisfied by their individual components. If even a single pair of components in the two mixtures is equivalent, the intersection of their supports creates a region of support overlap for both measures, precluding the existence of an error-free decision boundary. This underscores the importance of the condition $\gamma_{i}\perp\gamma_{j}\,\forall i\in I_{1},j\in I_{2}$ in high-dimensional settings.
@@ 5 Practical Applications
Since GMMs are universal approximators for continuous probability densities, the implications of Lemma 2.3 extend to nearly any dataset that can be modelled as a mixture of localised concentrations of measure. In high-dimensional anomaly detection, the theorem predicts that an anomaly, when modelled as a singular mixture component, will not disappear into the baseline data distribution as the dimensionality $d\to\infty$. Instead, the concentration of measure on thin shells ensures that these mixtures remain perfectly distinguishable, as their respective supports are disjoint in the infinite dimensional limit.
@@
{acks}
[Acknowledgments] The author would like to thank Prof. C. Schwab, that pointed out the Feldman-Hajék theorem to me, during a workshop on infinite dimensional statistics.

### v2

@@
[…]
@@ 1 Introduction
The study of probability measures in high-dimensional spaces is often dominated by the phenomenon of measure concentration [2, 3, 4, 5]. As one transitions to infinite-dimensional settings, the geometry of these measures changes radically, leading to the behavior described for Gaussian measures in 1958 by the Feldman-Hájek dichotomy [4, 5]. This result belongs to a broader class of dichotomy theorems for infinite-dimensional measures, the most notable being Kakutani’s theorem on infinite product measures [7]. The Feldman-Hajék theorem states that two Gaussian measures in an infinite-dimensional Hilbert space are either equivalent or mutually singular. While the behavior of individual Gaussian measures is well-documented, modern statistical learning frequently employs Gaussian Mixture Models (GMMs) to represent complex distributions, and thus it is a relevant questions under which conditions two different GMMs can be perfectly separable in infinite dimensions. Some general discussion about sums of random variables and mixtures can be found in the book by Vakhania et al. in [12], but not a specific result for GMMs.
@@
Several authors have studied GMM models in high dimensional settings, but nobody, to the best knowledge of the author, has given conditions for mutual singularity of GMM models in infinite dimensions. Almost all research work found dealt with the problem of classifying or learning GMM models, and were not interested in finding conditions for mutual singularity.
@@
It may be of interest to the reader to highlight some of the works related to learning GMMs. It should be emphasised that these contributions are in a certain sense connected to and conceptually aligned with the research presented in this work. In fact, if two GMMs are mutually singular, it also means that they are perfectly separable, or in other words, under the assumption that one finds the right classifier, one can classify them with zero missclassification error. A first interesting work in this direction, is the one by Mignacco et al. [9]. In their work, they studied the role of regularisation in a classification problem of high-dimensional two-mixture of noisy gaussians. Refinetti et al. studied the problem of classifying high-dimensional Gaussian mixtures, with a focus on neural networks [10] and showed the superiority of neural networks over kernel methods. Azizyan et al. in [1] also studied the problem of learning Gaussian mixtures with equal mixing weights in a high-dimensional settings under small-mean separation. Other works goes along the same line, in studying how to learn gaussian mixtures. The objective of this study is not to address that point directly, but rather to investigate how the Feldman–Hajék theorem can be generalized and extended to the setting of countable Gaussian mixture measures. As already mentioned, this work is important for the problem of learning mixtures, as it deals with the limit where error-free (perfect) classification becomes theoretically possible.
@@
The primary contribution of this work is the extension of the Feldman-Hájek theorem to countable Gaussian Mixture Models. Specifically, we establish the following. (i) We provide a rigorous proof that a countable mixture of Gaussian measures is a well-defined probability measure. (ii) We prove that the mutual singularity of all Gaussian components of the mixtures is a sufficient condition for the mutual singularity of the resulting mixtures. (iii) We discuss briefly the “mixed case” demonstrating how the existence of equivalent components breaks the dichotomy and leads to partial absolute continuity. By extending these results to GMMs, we provide the first steps for the theoretical foundation for the study of asymptotic separability in complex, multi-modal distributions.
@@
The rest of this paper is organised as follows. In Section 2, we establish the formal measure-theoretic framework for countable mixtures and prove that they are valid probability measures via Lemmas 2.1 and 2.3. Section 3 presents our main result, the Gaussian Mixture Dichotomy Theorem (Theorem 3.1), and its proof. Finally, Section 4 discusses the implications of these results.
@@ 2 Formalism and Gaussian Mixtures
Let $(X,\mathcal{B}(X))$ be a measurable space, where $\mathcal{B}(X)$ is a $\sigma$-algebra. A countable mixture is a probability measure $\mu$ defined for every $A\subseteq\mathcal{B}(X)$ as the following.
@@
where $\gamma_{k}$ are probability Gaussian measures on $(X,\mathcal{B}(X))$ characterised by the mean vector $m_{k}$ and by their covariance $\Sigma_{k}$. $\pi_{k}$ are mixing weights such that $\pi_{k}\geq 0$ and $\sum_{k=1}^{\infty}\pi_{k}=1$. In Lemma 2.3 we prove that $\mu$ itself is a probability measure. Lemma 2.1 is a stepping stone to prove lemma 2.3.
@@
Given a Gaussian measure $\gamma_{i}$, for any For any sequence of disjoint sets $\{A_{n}\}_{n=1}^{\infty}$ with $A_{n}\subset(X,\mathcal{B}(X))$ the sum $\sum_{k=1}^{\infty}\gamma_{i}(A_{k})$ exists and is finite.
Proof 2.2.
@@
But also by definition $\cup_{n=1}^{\infty}A_{n}\subseteq X$ and therefore
$\gamma_{i}\left(\bigcup_{n=1}^{\infty}A_{n}\right)\leq\gamma_{i}(X)=1$ (3)
Additionally, since each $\gamma_{i}(A_{n})\geq 0$ then the partial sums $s_{m}=\sum_{n=1}^{m}\gamma_{i}(A_{n})$ are monotonically increasing and bounded above by 1, hence the partial sums converge. This concludes the proof.
Lemma 2.3 ($\mu$ is a probability measure).
@@
where $\gamma_{k}$ are probability Gaussian measures on $(X,\mathcal{B}(X))$ characterised by the mean vector $m_{k}$ and by their covariance $\Sigma_{k}$. $\pi_{k}$ are mixing weights such that $\pi_{k}\geq 0$ and $\sum_{k=1}^{\infty}\pi_{k}=1$. $\mu$ is a probability measure.
Proof 2.4.
@@
Now let us consider two GMMs $\mu$ and $\nu$ defined over two disjoint sets of indices $I_{1}$ and $I_{2}$ similarly to Equation 1. Now we will prove the following lemma.
Lemma 2.5.
Consider two disjoint set of indeces $I_{1}$ and $I_{2}$. Given $\gamma_{i}\ i\in I_{1}$ and $\gamma_{j}\ j\in I_{2}$ gaussian measures such that $\gamma_{i}\perp\gamma_{j}$ (the two measures are mutually singular) $\forall i\in I_{1}$ and $\forall j\in I_{2}$. Defining
@@
With $\pi_{i}\geq 0$ and $\sum_{i=1}^{\infty}\pi_{i}=1$. it follows that $\mu\perp\nu$.
Proof 2.6.
To prove this lemma, we need to show that $\exists A$ such that $\mu(A)=1$ and $\nu(A)=0$. Let us construct such a set. Recall that, by assumption, $\gamma_{i}\perp\gamma_{j}$ (the two measures are mutually singular) $\forall i\in I_{1}$ and $\forall j\in I_{2}$, thus there exists a set $A_{ij}$ such that $\gamma_{i}(A_{ij})=1$ and $\gamma_{j}(A_{ij})=0$ $\forall i\in I_{1}$ and $\forall j\in I_{2}$. Consider the set
@@
$\bigcup_{i\in I_{1}}S_{i}\subseteq X\Rightarrow\mu\left(\bigcup_{i\in I_{1}}S_{i}\right)\leq\mu(X)=1$ (18)
@@
Or in other words we have proven that $1\leq\mu(A)\leq 1$ and therefore $\mu(A)=1$. Thus we have proven that $\mu\perp\nu$.
@@ 3 The Gaussian Mixture Dichotomy Theorem
Let $(X,\mathcal{B}(X))$ be a measurable space. Let $\mu$ and $\nu$ be two Gaussian Mixture Models (GMMs) defined over disjoint index sets $I_{1}$ and $I_{2}$ as
@@
    The mean difference $(m_{i}-m_{j})$ does not belong to the common Cameron-Martin Space $H=\Sigma_{i}^{1/2}(X)=\Sigma_{j}^{1/2}(X)$.
Proof 3.2.
By Lemma 2.3, if $\gamma_{i}\perp\gamma_{j}$ for all $i\in I_{1}$ and $j\in I_{2}$, then $\mu\perp\nu$. Under the standard Feldman-Hájek theorem, two Gaussian measures are mutually singular if and only if the aforementioned operator or mean conditions are violated. Since these conditions imply $\gamma_{i}\perp\gamma_{j}$, the singularity of the mixtures follows from the existence of the set $A=\bigcup_{i\in I_{1}}(\bigcap_{j\in I_{2}}A_{ij})$ such that $\mu(A)=1$ and $\nu(A)=0$.
@@ 4 Results and Discussion
In finite dimensions (e.g. $\mathbb{R}^{d}$), two non-degenerate Gaussian measures are almost always equivalent because their supports coincide with the entire space. However, singularity ($\gamma_{i}\perp\gamma_{j}$) can occur in finite dimensions if the measures are degenerate, for instance, or if they are supported on disjoint lower-dimensional affine subspaces.
The Gaussian Mixture Dichotomy Theorem becomes particularly relevant in infinite-dimensional spaces. In infinite-dimensional spaces, the dichotomy takes hold: two Gaussian measures are either equivalent or mutually singular. This rigorous disjointness of support in infinite dimensions is a characteristic property of Gaussian measures [4] and is in general valid for product-type measures [7] and, as we have demonstrated, scales to Gaussian Mixture Models (GMMs). If the components of the mixtures are singular, a condition almost always satisfied in the infinite-dimensional limit due to shifts in mean or scaling of covariance, the global mixtures $\mu$ and $\nu$ are themselves perfectly separable.
@@ 4.2 The Mixed Case
An important aspect in applying the Gaussian Mixture Dichotomy Theorem arises when the index sets $I_{1}$ and $I_{2}$ each include Gaussian pairs that are not all of the same type, but instead comprise a combination of pairs that are equivalent and pairs that are mutually singular. While the standard theorem for pure Gaussian measures permits only a binary state (equivalence or singularity), the mixture model $\mu$ and $\nu$ allows for a ”mixed” or intermediate state.
@@
In practical terms, this means that the strictly “all-or-nothing” nature of separability in GMMs fundamentally depends on the pair-wise singularity condition being satisfied by their individual components. If even a single pair of components in the two mixtures is equivalent, the intersection of their supports creates a region of support overlap for both measures, precluding the existence of an error-free decision boundary. This highlight the importance of the condition $\gamma_{i}\perp\gamma_{j}\,\forall i\in I_{1},j\in I_{2}$ in high-dimensional settings.
@@ 5 Practical Applications
Since GMMs are universal approximators for continuous probability densities, the implications of Lemma 2.5 extend to nearly any dataset that can be modelled as a mixture of localised concentrations of measure. In high-dimensional anomaly detection, the theorem predicts that an anomaly, when modelled as a singular mixture component, will not disappear into the baseline data distribution in the infinite-dimensional limit. Instead, the concentration of measure on thin shells ensures that these mixtures remain perfectly distinguishable, as their respective supports are disjoint in the infinite dimensional limit.